# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ A.3. The Fungible
The Fungible drop took place 12–14 April 2021 on Nifty Gateway, generating approximately \$16.8 million in sales over three days . The open edition “Cubes” were sold in timed windows at \$500, \$1,000, and \$1,500 price points on successive days, with a total of 23,598 cubes purchased. These merged into 6,156 NFTs , distributed to 1,659 unique buyers (reflecting many bought multiple cubes). Sotheby’s auctioned The Pixel for \$1.36 million and The Switch for \$1.44 million , each becoming one of the earliest NFT artworks to surpass \$1 million. The special one-of-one The Cube was awarded to the top cube buyer (who acquired 1,000 cubes), and 100 editions of Complexity were awarded to the next 100 top buyers . In total, over 3,080 transactions were recorded during the primary sale. At the time, it set a record for the largest NFT sale by a single artist on a platform, until Pak’s own Merge later that year.
@@
\$ASH is a social currency backed by curated extinction. The value, the utility and the identity of \$ASH is balanced by its users. \$ASH is a fair launch token that started with 0 supply. Liquidity Pools and Swaps are created by the community. (Uniswap) \$ASH can be mined by burning NFTs, bringing a dynamic balance to scarcity. For every NFT you burn you make the rest rarer, favoring their collectors. \$ASH currently has 2 burn tiers where Pak NFTs yield more ash than others. Both tiers follow the same bonding curve mechanism with different multipliers. Pak NFTs tier curve starts with 1000 \$ASH and others start with 2 \$ASH yield. Halving happens at 5M token supply impacting all burn tiers. You can burn a token by clicking the token image. You can also burn by providing an opensea link for the token.
Figure 5. Quotes from burn.art (2021)
@@
burn.art (and its native token \$ASH) is Pak’s poetic framework for “creation by destruction,” an artwork-platform that turns the act of burning NFTs into an artistic medium in itself. Launched in 2021 after The Fungible, it proposed a cyclical ecosystem: collectors are invited to destroy their existing NFT artworks (irreversibly sending them to a burn address), and in return they receive a new fungible token called ASH – essentially the “ashes” of the burned art (202, 2021b). These ashes (\$ASH) can then be spent on new art pieces by Pak, or even other creators, thereby continuing the cycle of creative destruction. The conceptual core is a commentary on value and permanence: by explicitly linking loss and creation, Pak challenges the notion of digital art as eternally replicable or static. Instead, burn.art frames destruction as a generative act, asking us to consider what we value more – the art we had, or the new possibilities unlocked by its sacrifice.
@@
The underlying mechanism of burn.art is an Ethereum smart contract that accepts NFTs and issues ERC-20 ASH tokens in return. Technically, when a user initiates a burn, the NFT is transferred to an inaccessible null address (removing it from circulation forever), while the contract credits the user with a certain amount of ASH according to a predefined schedule. The supply of ASH is thus dynamically generated by destruction: every token in circulation is proof that some piece of digital art was sacrificed. This introduces a self-balancing scarcity model – the more art is burned, the rarer the remaining artworks become, favoring their collectors by increasing relative scarcity. Pak designed \$ASH’s tokenomics such that all his future drops could interact with this system, for example by accepting only \$ASH as payment for certain new works or giving exclusive access to ASH holders. This created an evolving feedback loop: the dynamics of burning and creating continually affect each other, crafting a mini-economy where collectors’ decisions to destroy or hold collectively shape the progression of the art ecosystem.
@@
Using burn.art is itself a performative interaction: collectors must actively choose to irreversibly destroy one of their NFTs in order to partake. Upon connecting their crypto wallet to the burn.art interface, a user can select an NFT they own and send it to the burn contract; the system then “mines” (mints) a corresponding amount of ASH tokens as a reward. The amount of ASH received depends on the category of the burned NFT – Pak implemented a whitelist with multipliers, meaning more significant or scarce works yield more \$ASH. This process engages users on a psychological level: burning an artwork, especially a valuable one, is a moment of suspense and conviction, effectively making the user a co-creator of the conceptual piece. Communities sprung up around sharing burning experiences and strategies (e.g., which NFTs to burn for optimal \$ASH yield), turning what could have been a solitary act into a communal ritual. By requiring collectors to “prove” their dedication (through destruction) to gain access to new creations, burn.art ensured that participation was not passive but deeply intentional and symbolic.
@@
While burn.art introduced a new level of community agency (the choice of what and when to burn lies with the users), Pak maintained strategic control over the ecosystem’s parameters. He defined which NFTs are burnable for ASH and their relative yields via the whitelist — essentially curating the value of destruction. Furthermore, as the creator of the ASH token, Pak could influence its use by deciding which new artworks require ASH for purchase, thereby indirectly guiding demand for burning. That said, once the contracts were deployed, the day-to-day operation of the system is autonomous: any eligible NFT can be burned permissionlessly, and \$ASH is dispensed according to code, not at Pak’s discretion. In keeping with decentralization, the value of ASH and the decision of what to sacrifice for it were left to the open market and individual collectors. Pak’s control was thus front-loaded – in the design and rules of the game – but he stood back as participants carried out the dramatic act of creation-through-destruction, letting the social experiment unfold organically.
Collective Emerging Behavior
burn.art quickly garnered a devoted following of Pak’s collectors and intrigued onlookers, forming a sub-community fixated on the burn-to-mint ritual. Hundreds of NFTs were burned in the initial weeks, from common items to high-value artworks, as collectors sought to accumulate ASH either for status or to use in Pak’s future drops. This created a shared ethos among participants: a mix of fanaticism and camaraderie emerged, where those who burned valuable assets were celebrated (or playfully lamented) for their commitment to the art. Discussions on forums and social media centered on what NFTs people were willing to sacrifice; the “burn culture” introduced by Pak made some collectors view their holdings in a new light (knowing any Pak piece, if desired, could be converted to ASH, a kind of artistic afterlife for stagnant assets). Over time, the collective action of these participants literally shaped Pak’s art market: as more tokens were destroyed, the remaining ones grew rarer, and owning ASH became a badge of participation in Pak’s evolving narrative. This self-selected community, willing to destroy to create, exemplified a new kind of collector ethos unique to the protocol – one that values conceptual engagement over simply accumulating objects.
@@
burn.art and ASH elevate an almost mythic cycle of death and rebirth within the digital realm, making it one of Pak’s most overtly metaphorical works. The project transforms the destructive act of blockchain burning into a meaningful ritual, suggesting that even in a space of infinite reproducibility, sacrifice can carry weight and give rise to new beauty. The poetry here lies in how absence is turned into presence: each ASH token is a memorial of a destroyed artwork and simultaneously the seed of a new creation. In a world where digital files ostensibly last forever, Pak introduced impermanence and choice as artistic elements – echoing the destruction of physical art (recalling Gustav Metzger’s “auto-destructive art” or Jean Tinguely’s self-destroying sculptures) but with a constructive twist. The broader impact of burn.art is its commentary on value: it forces the community to question what gives an artwork value – its continued existence, or the legacy it leaves through transformation. By making his audience perform a collective dance of loss and renewal, Pak illustrated that in digital ecosystems, as in nature, endings can be beginnings. This piece thus bridges conceptual art and social experiment, leaving a lasting impression about the responsibilities and powers that protocols confer upon both artists and collectors.
@@
burn.art launched in April 2021. Within the first 48 hours, users burned over 1,000 NFTs to generate \$ASH. By February 2022, more than 20,000 NFTs had been burned via the platform , including high-profile pieces by Pak and other artists, yielding over 4 million \$ASH tokens. \$ASH reached a peak market capitalization above \$100 million during the 2021 NFT boom, reflecting the community’s speculative interest. Pak’s first \$ASH-only NFT drop, “Chapter One – Carbon,” sold out in minutes, consuming a substantial portion of the token supply. The initial reward rate was 1000 \$ASH for Pak NFTs and 2 \$ASH for others , with a halving that occurred in mid-2022 reducing outputs by half. As of 2025, over 30 million \$ASH have been minted through successive burns, and the token is used by a handful of other crypto artists who accept \$ASH for their work, extending Pak’s poetic economy of art and ash.
@@ A.5. Lost Poets
Figure 6. Lost Poets
Figure 7. Lost Poets after revealing
@@
Lost Poets (September 2021) is a large-scale NFT art project that merges generative art, narrative puzzle, and strategic game into one poetic whole. At its core, the project consists of 65,536 unique AI-generated portrait NFTs (“Poets”) originating from 1,024 rarer “Origin” Poets (Anderson, [n. d.]), all imbued with an aura of mystery and antiquity. The concept draws inspiration from Borges’s Library of Babel and the idea of rediscovering lost creative souls: Pak envisioned these Poets as ancient voices “not found” until the community brings them to light. The project was unveiled as a multi-act saga – a kind of unfolding story where the NFTs would evolve over time – thereby exploring themes of memory (lost knowledge rediscovered), evolution (the NFTs reveal attributes in phases), and collective intelligence. Lost Poets is fundamentally about the interplay between algorithmic creation (the AI art and smart-contract logic) and human participation in meaning-making: it asks what happens when thousands of collectors jointly play a literary game with art, time, and chance.
@@
The Lost Poets smart contracts orchestrated an intricate multi-stage evolution of the NFTs. Initially, 65,536 Page tokens were minted (with Pak reserving 1,024 for the special Origin Poets). These ERC-1155 tokens had built-in utility: a holder could “burn” a Page token to transform it into a Poet token once Act II commenced . The transformation was irreversible, effectively migrating the NFT from one state (Page) to another (Poet) and revealing the AI-generated portrait and some initial attributes. Crucially, the contract did not reveal all attributes of a Poet at once; instead, additional traits (like words the Poet could “speak”) unlocked gradually over Act II and Act III, either automatically or through user actions. Feeding a Page to a Poet (another burn action) was a method to update the Poet’s metadata – the contract would record the new name given by the user and randomly assign a set of words to that Poet, altering its uniqueness. All these state changes were governed by on-chain logic, ensuring that no two Poets ended up the same despite coming from identical Pages. The system also tracked leaderboards and triggers for awarding the Origin Poets to top collectors at the end of Act I. Finally, a time-lock dynamic was present: at the conclusion of the one-year project timeline, the contract “locked in” the final state of all Poets (Act IV, aptly titled “The Twist”) and dispensed promised ASH rewards to participants, marking the end of the journey. This layered smart contract design allowed Lost Poets to behave almost like an autonomous game engine running on Ethereum, driving the content and rarity of the NFTs in response to player inputs and time.
@@
Engagement with Lost Poets was deeply interactive and unfolded in distinct phases (Acts I–IV). It began with the distribution of “Pages” – blank NFTs that served as tickets to the game (Los, 2023). Early on, Pak rewarded loyal fans by airdropping Pages to wallets holding at least 25 ASH tokens and opened a public sale where others could purchase Pages (priced at 0.32 ETH each). Once armed with Pages, participants faced choices in Act II (“The Reveal”): they could burn a Page token to summon its Poet (revealing a unique AI-crafted portrait), or they could even burn Pages for more ASH instead, adding a twist of sacrifice to the game. In Act III (“The Explorer”), collectors could feed extra Pages to their Poets, each feed allowing them to rename the Poet and gifting the character new words (2–4 random words per Page) which enriched its traits and lore. Throughout these interactions, Pak maintained a live leaderboard; the top 100 collectors who amassed the most Pages received additional rare Origin Poets as rewards, incentivizing strategic play and competition. The project was designed to span 365 days, during which participants continuously engaged – naming their Poets, deciphering clues, trading Pages and Poets, and anticipating “The Twist” of Act IV. This sustained, game-like interaction meant that collecting Lost Poets was not a one-off transaction but a prolonged creative exercise that bound the community together.
@@
Pak’s role in Lost Poets was akin to a dungeon master setting the rules of a game, then letting the players roam. He exercised significant control in the design phase: determining the number of tokens, the pacing of the acts, the mechanics for conversion and feeding, and even withholding the 1,024 Origin Poets to distribute as rewards or for future surprises. However, once the project commenced, the progression was largely automated and community-driven. Pak did not alter the contract rules mid-course; the code itself dictated when new attributes would appear and how each action translated into an outcome. The element of the unknown (“The Twist” of Act IV) was built-in – Pak had likely pre-planned an ending but kept it secret, triggering it through the programmed schedule or an on-chain call at the appropriate time. Throughout the year-long evolution, Pak’s direct intervention was minimal, aside from providing occasional hints or narrative flavor via social media to enrich the lore. In essence, after launching Lost Poets, he relinquished control to the protocol and the participants, allowing the artwork to self-evolve within the boundaries he had coded. This balance – tight authorial control over structure, but freedom for users within it – underscores Pak’s commitment to exploring protocol as an art form.
Collective Emerging Behavior
Lost Poets attracted a massive and engaged community, as evidenced by the entire supply of 65,536 Pages selling out in just 2 hours (raising about \$70 million) and thousands of collectors joining the fray. The project’s gamified nature led to rich emergent behaviors: collectors formed online groups to decode hidden clues in Poet attributes and to strategize the best use of Pages. A vibrant secondary market developed where Pages and Poets were traded, with some speculators hoarding Pages early in hopes of obtaining more Origins or leveraging them in later acts. The competition for the top 100 collector spots was intense, leading certain individuals to accumulate enormous quantities of Pages (and become known figures in the community for it). Meanwhile, other participants took a more curatorial approach, carefully naming their Poets and sharing the whimsical or profound word combinations their Poets “spoke” after feeding – effectively collaboratively writing a decentralized poem through their NFTs. Each phase change (Act II’s reveal, Act III’s feeding, etc.) was accompanied by collective excitement on forums and Discord, as people shared discoveries and theories about what Act IV, “The Twist,” might entail. By the end of the 365 days, Lost Poets had fostered not just a market but a participatory culture – blending competition, storytelling, and collaboration – arguably one of the most sustained communal engagements for an NFT project of its era.
@@
The poetry of Lost Poets emerges through its synthesis of technology and storytelling. On one level, it is a meditation on authorship: each collector who names their Poet and contributes words becomes a co-author, blurring the line between artist and audience in a manner reminiscent of exquisite corpse or collaborative poetry. The project also poignantly plays with the idea of “lost” voices found – the algorithm conjured tens of thousands of unique faces and fragments of text, like ghosts of poets past, and gave the community the power to nourish these ghosts with new words. Over the course of the acts, participants experienced themes of discovery, transformation, and ephemerality (as unused Pages dwindled and choices had to be made). In the final Epilogue, when all Poets reached their ultimate form and the last secrets were unveiled, the community could look back on a journey that it had collectively authored. Pak managed to reveal that NFTs need not be static collectibles; they can be alive with narrative potential and participant-driven change. Lost Poets stands as a testament to the idea that protocol-based art can achieve a form of literary and artistic richness – it turned a blockchain ledger into the stage for a year-long participatory saga about remembering and reinventing creative voices. The broader impact is an expanded notion of what digital art can be: not just an image or an object, but a living story that engages its audience in poetic meaning-making.

### v2

@@ A.3. The Fungible
The Fungible drop took place 12–14 April 2021 on Nifty Gateway, generating approximately \$16.8 million in sales over three days. The open edition “Cubes” were sold in timed windows at \$500, \$1,000, and \$1,500 price points on successive days, with a total of 23,598 cubes purchased. These merged into 6,165 NFTs, distributed to 3,080 unique buyers (reflecting many bought multiple cubes). Sotheby’s auctioned The Pixel for \$1.36 million and The Switch for \$1.44 million, each becoming one of the earliest NFT artworks to surpass \$1 million. The special one-of-one The Cube was awarded to the top cube buyer, and 100 editions of Complexity were awarded to the next 100 top buyers. In total, over 3,080 unique buyers participated in the primary sale. At the time, it set a record for the largest NFT sale by a single artist on a platform, until Pak’s own Merge later that year.
Annotation
The Fungible is the portfolio’s strongest early case for C5 (economy-driven engagement): the Sotheby’s partnership, tiered pricing, leaderboard competitions, and puzzle rewards made the market event indistinguishable from the artwork. It equally demonstrates C1 (system-centric composition), as the algorithmic conversion of fungible cubes into unique NFTs—whereby participants could purchase interchangeable units and watch an algorithm transmute them into bespoke non-fungible artworks—embodies the shift from object to process. Crucially, the work plays with the very boundary between fungible and non-fungible: collectors could combine non-fungible elements back into fungible form, a kind of “coin play” that literalizes the conceptual slippage between currency and art. C3 (distributed agency) is present but structured—participants co-determine the outcome through purchasing decisions, yet within tightly scripted rules, distinguishing this from the more open-ended authorship of Lost Poets or Censored. What elevates The Fungible beyond a single project, however, is its role as the foundational experiment in composability for Pak’s subsequent practice. The market here provides an interaction space in which users can play, combine elements, and test composability, and the mechanics piloted in The Fungible directly laid the groundwork for the token-burning logic of burn.art, the multi-stage transformations of Lost Poets, and the self-merging dynamics of Merge. In this respect, The Fungible constitutes an early-stage laboratory for C7 (interoperability and composability), leading users toward open-ended authorship and establishing the paradigm that Pak would elaborate across his entire oeuvre.
@@
Refer to caption A text-based screenshot of the burn.art website explaining the \$ASH token system. The page describes \$ASH as a social currency backed by curated extinction, explains that it can be mined by burning NFTs, details a two-tier system where Pak NFTs yield more \$ASH, mentions a bonding mechanism with different multipliers, a halving event at 1,000 \$ASH and another at 5 million token supply, and provides instructions for adding \$ASH to MetaMask.
Figure 5. The burn.art landing page (2021), outlining the \$ASH token economy. The system formalizes “creation through destruction”: collectors permanently burn NFTs to receive \$ASH tokens, establishing a cyclical protocol where artistic value is destroyed, transmuted, and regenerated.
@@
burn.art (and its native token \$ASH) is Pak’s poetic framework for “creation by destruction,” an artwork-platform that turns the act of burning NFTs into an artistic medium in itself. Launched in 2021 after The Fungible, it proposed a cyclical ecosystem: collectors are invited to destroy their existing NFT artworks (irreversibly sending them to a burn address), and in return they receive a new fungible token called ASH—essentially the “ashes” of the burned art (.ART, 2021). These ashes (\$ASH) can then be spent on new art pieces by Pak, or even other creators, thereby continuing the cycle of creative destruction. The conceptual core is a commentary on value and permanence: by explicitly linking loss and creation, Pak challenges the notion of digital art as eternally replicable or static. Instead, burn.art frames destruction as a generative act, asking us to consider what we value more—the art we had, or the new possibilities unlocked by its sacrifice.
@@
The underlying mechanism of burn.art is an Ethereum smart contract that accepts NFTs and issues ERC-20 ASH tokens in return. Technically, when a user initiates a burn, the NFT is transferred to an inaccessible null address (removing it from circulation forever), while the contract credits the user with a certain amount of ASH according to a predefined schedule. The supply of ASH is thus dynamically generated by destruction: every token in circulation is proof that some piece of digital art was sacrificed. This introduces a self-balancing scarcity model—the more art is burned, the rarer the remaining artworks become, favoring their collectors by increasing relative scarcity. Pak designed \$ASH’s tokenomics such that all his future drops could interact with this system, for example by accepting only \$ASH as payment for certain new works or giving exclusive access to ASH holders. This created an evolving feedback loop: the dynamics of burning and creating continually affect each other, crafting a mini-economy where collectors’ decisions to destroy or hold collectively shape the progression of the art ecosystem.
@@
Using burn.art is itself a performative interaction: collectors must actively choose to irreversibly destroy one of their NFTs in order to partake. Upon connecting their crypto wallet to the burn.art interface, a user can select an NFT they own and send it to the burn contract; the system then “mines” (mints) a corresponding amount of ASH tokens as a reward. The amount of ASH received depends on the category of the burned NFT—Pak implemented a whitelist with multipliers, meaning more significant or scarce works yield more \$ASH. This process engages users on a psychological level: burning an artwork, especially a valuable one, is a moment of suspense and conviction, effectively making the user a co-creator of the conceptual piece. Communities sprung up around sharing burning experiences and strategies (e.g., which NFTs to burn for optimal \$ASH yield), turning what could have been a solitary act into a communal ritual. By requiring collectors to “prove” their dedication (through destruction) to gain access to new creations, burn.art ensured that participation was not passive but deeply intentional and symbolic.
@@
While burn.art introduced a new level of community agency (the choice of what and when to burn lies with the users), Pak maintained strategic control over the ecosystem’s parameters. He defined which NFTs are burnable for ASH and their relative yields via the whitelist—essentially curating the value of destruction. Furthermore, as the creator of the ASH token, Pak could influence its use by deciding which new artworks require ASH for purchase, thereby indirectly guiding demand for burning. That said, once the contracts were deployed, the day-to-day operation of the system is autonomous: any eligible NFT can be burned permissionlessly, and \$ASH is dispensed according to code, not at Pak’s discretion. In keeping with decentralization, the value of ASH and the decision of what to sacrifice for it were left to the open market and individual collectors. Pak’s control was thus front-loaded—in the design and rules of the game—but he stood back as participants carried out the dramatic act of creation-through-destruction, letting the social experiment unfold organically.
Collective Emergent Behavior
burn.art quickly garnered a devoted following of Pak’s collectors and intrigued onlookers, forming a sub-community fixated on the burn-to-mint ritual. Hundreds of NFTs were burned in the initial weeks, from common items to high-value artworks, as collectors sought to accumulate \$ASH either for status or to use in Pak’s future drops. This created a shared ethos among participants: a mix of fanaticism and camaraderie emerged, where those who burned valuable assets were celebrated (or playfully lamented) for their commitment to the art. Discussions on forums and social media centered on what NFTs people were willing to sacrifice; the “burn culture” introduced by Pak made some collectors view their holdings in a new light (knowing any Pak piece, if desired, could be converted to \$ASH, a kind of artistic afterlife for stagnant assets). Over time, the collective action of these participants literally shaped Pak’s art market: as more tokens were destroyed, the remaining ones grew rarer, and owning \$ASH became a badge of participation in Pak’s evolving narrative. This self-selected community, willing to destroy to create, exemplified a new kind of collector ethos unique to the protocol—one that values conceptual engagement over simply accumulating objects.
@@
burn.art and \$ASH elevate an almost mythic cycle of death and rebirth within the digital realm, making it one of Pak’s most overtly metaphorical works. The project transforms the destructive act of blockchain burning into a meaningful ritual, suggesting that even in a space of infinite reproducibility, sacrifice can carry weight and give rise to new beauty. The poetry here lies in how absence is turned into presence: each \$ASH token is a memorial of a destroyed artwork and simultaneously the seed of a new creation. In a world where digital files ostensibly last forever, Pak introduced impermanence and choice as artistic elements—echoing the destruction of physical art (recalling Gustav Metzger’s “auto-destructive art” or Jean Tinguely’s self-destroying sculptures) but with a constructive twist. The broader impact of burn.art is its commentary on value: it forces the community to question what gives an artwork value—its continued existence, or the legacy it leaves through transformation. By making his audience perform a collective dance of loss and renewal, Pak illustrated that in digital ecosystems, as in nature, endings can be beginnings. This piece thus bridges conceptual art and social experiment, leaving a lasting impression about the responsibilities and powers that protocols confer upon both artists and collectors.
@@
burn.art launched in May 2021. Within the first 48 hours, users burned over 1,000 NFTs to generate \$ASH. By February 2022, more than 20,000 NFTs had been burned via the platform, including high-profile pieces by Pak and other artists, yielding over 4 million \$ASH tokens. \$ASH reached a peak market capitalization above \$100 million during the 2021 NFT boom, reflecting the community’s speculative interest. Pak’s first \$ASH-only NFT drop, “Chapter One—Carbon,” sold out in minutes, consuming a substantial portion of the token supply. The initial reward rate was 1000 \$ASH for Pak NFTs and 2 \$ASH for others, with an automatic halving mechanism that reduces rewards at each supply milestone. As of 2025, over 30 million \$ASH have been minted through successive burns, and the token is used by a handful of other crypto artists who accept \$ASH for their work, extending Pak’s poetic economy of art and \$ASH.
Annotation
burn.art is the pivotal work for C7 (interoperability and composability): it introduces \$ASH as a cross-project fungible token that bridges The Fungible, Lost Poets, Merge, and future drops, transforming Pak’s oeuvre from a series of discrete works into a unified economic ecosystem—the center of the Pak universe. It is also a powerful case for C6 (poetic message embedding), as the ritual of destruction-as-creation literalizes themes of sacrifice, renewal, and the cyclical nature of value. C2 (autonomous governance) is strong—the burn contract operates permissionlessly—yet Pak retains unusual ongoing influence through an ultimate DAO behind the contract that can tune burning rates for each series, granting dynamic ability to control floor prices across the entire collection. This mechanism functions as an economic adjustment machine: if the floor price of any Pak series drops below the potential \$ASH yield from burning, buyers have a rational incentive to purchase and burn those NFTs into \$ASH, thereby automatically establishing a guaranteed floor price. The structure operates, in effect, like a Federal Reserve for Pak’s own art—a brilliantly engineered economic mechanism that guarantees basic value while preserving market freedom. This hybrid of autonomy and curatorial control complicates any simple narrative of decentralization in protocol art, making burn.art the only project where the artist actively steers post-launch demand through parametric governance rather than direct intervention. In its synthesis of permissionless burning, cross-project interoperability, and dynamic economic tuning, burn.art is arguably the most ingenious infrastructure-level artwork in the portfolio: an incredibly smart mechanism that turns the entire ecosystem into a self-regulating economy of creative destruction.
@@ A.5. Lost Poets
Refer to caption Two NFT listings side by side on a marketplace. Left: a blank gray square labeled “Page” priced at 0.0078 ETH, representing an unrevealed token. Right: an AI-generated portrait labeled number 6365 priced at 0.0449 ETH, showing a Renaissance-style female face with mechanical elements, representing a revealed Poet.
Figure 6. Lost Poets: before and after the reveal. Left: an unrevealed Page token, a blank placeholder. Right: a revealed Poet (number 6365), an AI-generated portrait unlocked by burning the Page—illustrating the irreversible state transition at the heart of the project’s multi-act protocol.
A wide screenshot of an OpenSea-style gallery showing approximately thirty revealed Lost Poets NFTs in a grid. Each displays a unique AI-generated portrait in diverse styles—some photorealistic, some painterly, some surreal—depicting faces of varying ages, ethnicities, and visual treatments. A sidebar on the left shows filtering options for collection attributes.
Figure 7. Gallery view of revealed Lost Poets on the secondary market. Each portrait was procedurally generated with unique visual traits; collectors further individualized their Poets by naming them and “feeding” additional Page tokens to unlock poetic word combinations, layering human authorship onto algorithmic creation.
@@
Lost Poets (September 2021) is a large-scale NFT art project that merges generative art, narrative puzzle, and strategic game into one poetic whole. At its core, the project consists of 65,536 unique AI-generated portrait NFTs (“Poets”) originating from 1,024 rarer “Origin” Poets (Anderson, 2022), all imbued with an aura of mystery and antiquity. The concept draws inspiration from Borges’s Library of Babel and the idea of rediscovering lost creative souls: Pak envisioned these Poets as ancient voices “not found” until the community brings them to light. The project was unveiled as a multi-act saga—a kind of unfolding story where the NFTs would evolve over time—thereby exploring themes of memory (lost knowledge rediscovered), evolution (the NFTs reveal attributes in phases), and collective intelligence. Lost Poets is fundamentally about the interplay between algorithmic creation (the AI art and smart-contract logic) and human participation in meaning-making: it asks what happens when thousands of collectors jointly play a literary game with art, time, and chance.
@@
The Lost Poets smart contracts orchestrated an intricate multi-stage evolution of the NFTs. Initially, 65,536 Page tokens were minted (with Pak reserving 1,024 for the special Origin Poets). These ERC-1155 tokens had built-in utility: a holder could “burn” a Page token to transform it into a Poet token once Act II commenced. The transformation was irreversible, effectively migrating the NFT from one state (Page) to another (Poet) and revealing the AI-generated portrait and some initial attributes. Crucially, the contract did not reveal all attributes of a Poet at once; instead, additional traits (like words the Poet could “speak”) unlocked gradually over Act II and Act III, either automatically or through user actions. Feeding a Page to a Poet (another burn action) was a method to update the Poet’s metadata—the contract would record the new name given by the user and randomly assign a set of words to that Poet, altering its uniqueness. All these state changes were governed by on-chain logic, ensuring that no two Poets ended up the same despite coming from identical Pages. The system also tracked leaderboards and triggers for awarding the Origin Poets to top collectors at the end of Act I. Finally, a time-lock dynamic was present: at the conclusion of the one-year project timeline, the contract “locked in” the final state of all Poets (Act IV, aptly titled “The Twist”) and dispensed promised \$ASH rewards to participants, marking the end of the journey. This layered smart contract design allowed Lost Poets to behave almost like an autonomous game engine running on Ethereum, driving the content and rarity of the NFTs in response to player inputs and time.
@@
Engagement with Lost Poets was deeply interactive and unfolded in distinct phases (Acts I–IV). It began with the distribution of “Pages”—blank NFTs that served as tickets to the game (NFT Metria, 2023). Early on, Pak rewarded loyal fans by airdropping Pages to wallets holding at least 25 \$ASH tokens and opened a public sale where others could purchase Pages (priced at 0.32 ETH each). Once armed with Pages, participants faced choices in Act II (“The Reveal”): they could burn a Page token to summon its Poet (revealing a unique AI-crafted portrait), or they could even burn Pages for more \$ASH instead, adding a twist of sacrifice to the game. In Act III (“The Explorer”), collectors could feed extra Pages to their Poets, each feed allowing them to rename the Poet and gifting the character new words (2–4 random words per Page) which enriched its traits and lore. Throughout these interactions, Pak maintained a live leaderboard; the top 100 collectors who amassed the most Pages received additional rare Origin Poets as rewards, incentivizing strategic play and competition. The project was designed to span 365 days, during which participants continuously engaged—naming their Poets, deciphering clues, trading Pages and Poets, and anticipating “The Twist” of Act IV. This sustained, game-like interaction meant that collecting Lost Poets was not a one-off transaction but a prolonged creative exercise that bound the community together.
@@
Pak’s role in Lost Poets was akin to a dungeon master setting the rules of a game, then letting the players roam. Pak exercised significant control in the design phase: determining the number of tokens, the pacing of the acts, the mechanics for conversion and feeding, and even withholding the 1,024 Origin Poets to distribute as rewards or for future surprises. However, once the project commenced, the progression was largely automated and community-driven. Pak did not alter the contract rules mid-course; the code itself dictated when new attributes would appear and how each action translated into an outcome. The element of the unknown (“The Twist” of Act IV) was built-in—Pak had likely pre-planned an ending but kept it secret, triggering it through the programmed schedule or an on-chain call at the appropriate time. Throughout the year-long evolution, Pak’s direct intervention was minimal, aside from providing occasional hints or narrative flavor via social media to enrich the lore. In essence, after launching Lost Poets, Pak relinquished control to the protocol and the participants, allowing the artwork to self-evolve within the boundaries Pak had coded. This balance—tight authorial control over structure, but freedom for users within it—underscores Pak’s commitment to exploring protocol as an art form.
Collective Emergent Behavior
Lost Poets attracted a massive and engaged community, as evidenced by the entire supply of 65,536 Pages selling out in just 2 hours (raising about \$70 million) and thousands of collectors joining the fray. The project’s gamified nature led to rich emergent behaviors: collectors formed online groups to decode hidden clues in Poet attributes and to strategize the best use of Pages. A vibrant secondary market developed where Pages and Poets were traded, with some speculators hoarding Pages early in hopes of obtaining more Origins or leveraging them in later acts. The competition for the top 100 collector spots was intense, leading certain individuals to accumulate enormous quantities of Pages (and become known figures in the community for it). Meanwhile, other participants took a more curatorial approach, carefully naming their Poets and sharing the whimsical or profound word combinations their Poets “spoke” after feeding—effectively collaboratively writing a decentralized poem through their NFTs. Each phase change (Act II’s reveal, Act III’s feeding, etc.) was accompanied by collective excitement on forums and Discord, as people shared discoveries and theories about what Act IV, “The Twist,” might entail. By the end of the 365 days, Lost Poets had fostered not just a market but a participatory culture—blending competition, storytelling, and collaboration—arguably one of the most sustained communal engagements for an NFT project of its era.
@@
The poetry of Lost Poets emerges through its synthesis of technology and storytelling. On one level, it is a meditation on authorship: each collector who names their Poet and contributes words becomes a co-author, blurring the line between artist and audience in a manner reminiscent of exquisite corpse or collaborative poetry. The project also poignantly plays with the idea of “lost” voices found—the algorithm conjured tens of thousands of unique faces and fragments of text, like ghosts of poets past, and gave the community the power to nourish these ghosts with new words. Over the course of the acts, participants experienced themes of discovery, transformation, and ephemerality (as unused Pages dwindled and choices had to be made). In the final Epilogue, when all Poets reached their ultimate form and the last secrets were unveiled, the community could look back on a journey that it had collectively authored. Pak managed to reveal that NFTs need not be static collectibles; they can be alive with narrative potential and participant-driven change. Lost Poets stands as a testament to the idea that protocol-based art can achieve a form of literary and artistic richness—it turned a blockchain ledger into the stage for a year-long participatory saga about remembering and reinventing creative voices. The broader impact is an expanded notion of what digital art can be: not just an image or an object, but a living story that engages its audience in poetic meaning-making.